OpenStudy (anonymous):

solve for real numbers x and y: (2y-3x) + (3x+4)i =1+ i

zepdrix (zepdrix):

\[\Large\bf\sf (2y-3x) + (3x+4)\mathcal i =1+ \mathcal i\]For this equation to hold true, the real part must equal the real part, and the imaginary parts must be equal as well.\[\Large\bf\sf 1\quad=\quad 2y-3x\]\[\Large\bf\sf \mathcal i\quad=\quad (3x+4)\mathcal i\]

zepdrix (zepdrix):

Look at the second equation, do you see how to solve for x?

OpenStudy (anonymous):

would I distribute the i to be 3xi-4i?

zepdrix (zepdrix):

No. We want to find real values for x. So we should probably try to find a way to get rid of the i's instead of distributing them.

zepdrix (zepdrix):

The i is multiplying those brackets. Let's do the inverse of that to get rid of the i's.

zepdrix (zepdrix):

Start by dividing each side by i.

OpenStudy (anonymous):

thats for the 3x+4i=i right?

zepdrix (zepdrix):

ya.

OpenStudy (anonymous):

would that be 3x+4=i??

zepdrix (zepdrix):

No, the i's divide out of each side.\[\Large\bf\sf \frac{\cancel{\mathcal i}}{\cancel{\mathcal i}}\quad=\quad \frac{(3x+4)\cancel{\mathcal i}}{\cancel{\mathcal i}}\]

zepdrix (zepdrix):

\[\Large\bf\sf 1\quad=\quad 3x+4\]

OpenStudy (anonymous):

oh I see so now I need to solve for both equations separately?

zepdrix (zepdrix):

Yes. This equation will allow you to solve for x. Then after you have x, you can plug it into the other equation to find your y.

OpenStudy (anonymous):

oh so I solved 1=3x+4 and I got -1 so I plug that into 2y-3x=1 to get y?

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

Thank you so much! :)